# A Theory of Learning with Autoregressive Chain of Thought

### v1

@@ 1 Introduction
In our view, a central component of autoregressive generation is that it is time-invariant, i.e., the same next-token-generator, with the identical parameters (e.g. a transformer with the same weights) is used at each step of generation, to generate each token as a function of the prefix thus far. Throughout the paper we emphasize how such time-invariance is crucial for allowing learning with sample complexity independent (perhaps up to a log-factor) of the generation length, and later on, of the compute time (number of steps) of a learned process. In this crucial regard, we deviate significantly from another recent (and inspiring) attempt to formalize autoregressive learning by Malach (2023), who considered time-dependent autoregressive learning. We directly contrast with Malach’s approach as we progress through our presentation. We begin by presenting a formal setting for time-invariant autoregressive Chain-of-Thought generation and learning using an abstract base class of next-token-generators (Section 2), studying the basics of learning in this framework for general and abstract base classes (Section 3), and, as an example, for what is perhaps the simplest possible base class, namely linear thresholds (Section 4). Already here we can start seeing the computational benefits of Chain-of-Thought learning and the statistical benefits of time-invariant autoregressive generation. We then turn to the more ambitious goal of “universal” learning of computable functions and see how far the benefits of time-invariance and autoregressive Chain-of-Thought take us, and how attention arises naturally (Sections 5 and 6).
@@ 2 Time-Invariant Autoregressive Chain-of-Thought Learning
For a finite set of tokens $\Sigma$, a next token generator is a mapping $f:\Sigma^{*}\to\Sigma$. On input ${\bm{x}}\in\Sigma^{*}$, autoregressive Chain-of-Thought (CoT) generation will start with a string containing only the input and will repeatedly apply $f$ to the string and append the newly generated token to it (see Figure 1). Formally, for a next-token generator $f$, we define the apply-and-append mapping $\overline{f}({\bm{x}}):{\bm{x}}\mapsto\textnormal{append}({\bm{x}},f({\bm{x}}))\in(\Sigma^{\left\lvert{{\bm{x}}}\right\rvert+1}$), [^2] For any ${\bm{x}}\in\Sigma^{*}$, we use $|{\bm{x}}|$ to denote its length. We use negative indices to refer to elements of a vector or string from the end, i.e. ${\bm{v}}[-1]$ is the last element. We use inclusive slice indexing, so that ${\bm{v}}[i:j]$ are the elements from position $i$ to $j$ inclusive. Dropping $i$ or $j$ means that the slice extends from the beginning or to the end, respectively. and apply $\overline{f}$ iteratively for $T$ steps to obtain a mapping from ${\bm{x}}$ to its $T$-step Chain-of-Thought:
@@
For any base class =${\mathcal{F}}=\{f:\Sigma^{*}\rightarrow\Sigma\}\subseteq\Sigma^{\Sigma^{*}}$ of generators, and a generation length $T$, we define the corresponding end-to-end and CoT classes:
@@
For simplicity, throughout the paper we consider a fixed thought generation length (i.e. number of iterations) $T$. This number of iterations $T$ will be a crucial parameter for us and we will study how the learning complexity and expressivity depend on $T$333 One could also define a variant with variable thought lengths and, e.g., a special token indicating when to stop. But if we would anyway be interested in the maximal thought length, we could just as well extend each thought with a padding token to make them all of the same length.
@@
For a base class ${\mathcal{F}}\subseteq\Sigma^{\Sigma^{*}}$, and generation length $T\in{\mathbb{N}}_{+}$, we say that ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf e2e}$-learnable over a domain ${\mathcal{X}}\subseteq\Sigma^{*}$, with sample complexity $m(\varepsilon,\delta)$ if there exists a learning rule $A:({\mathcal{X}}\times\Sigma)^{*}\to\Sigma^{{\mathcal{X}}}$ such that for every distribution ${\mathcal{D}}$ over ${\mathcal{X}}$ and $f_{*}\in{\mathcal{F}}$, for every $10<\varepsilon,\delta<$ 1, with probability at least $1-\delta$ over ${\bm{x}}_{1},\dots,{\bm{x}}_{m}\sim_{iid}{\mathcal{D}}$, with $S_{{\sf e2e}}=(({\bm{x}}_{1},y_{1}),\dots,({\bm{x}}_{m},y_{m}))$, $y_{i}=f_{*}^{{\sf e2e}\textnormal{-}{T}}({\bm{x}}_{i})$ and $m=m(\varepsilon,\delta)$, we have $L_{{\mathcal{D}},f_{*}}^{\textnormal{0-1}}(A(S_{{\sf e2e}}))\leq\varepsilon$.
Definition 1 amounts precisely to the standard definition of PAC-learning the hypothesis class ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$, but we give it explicitly in order to contrast it with learning in situations where the entire Chain-of-Thought is available during training:
@@
For a base class ${\mathcal{F}}\subseteq\Sigma^{\Sigma^{*}}$ and generation length $T\in{\mathbb{N}}_{+}$, we say that [^4] We emphasize that ${\sf CoT}$-learnability is a property of the base class ${\mathcal{F}}$ and length $T$, not only of the set ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$. ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf CoT}$-learnable over a domain ${\mathcal{X}}\subseteq\Sigma^{*}$, with sample complexity $m(\varepsilon,\delta)$, if there exists a learning rule $A:({\mathcal{X}}\times\Sigma^{T})^{*}\to\Sigma^{{\mathcal{X}}}$ such that for every distribution ${\mathcal{D}}$ over ${\mathcal{X}}$ and $f_{*}\in{\mathcal{F}}$, for every $10<\varepsilon,\delta<$ 1, with probability at least $1-\delta$ over ${\bm{x}}_{1},\dots,{\bm{x}}_{m}\sim_{iid}{\mathcal{D}}$, with $S_{{\sf CoT}}=({\bm{z}}_{1},\dots,{\bm{z}}_{m})$ where ${\bm{z}}_{i}=f_{*}^{{\sf CoT}\textnormal{-}{T}}({\bm{x}}_{i})$, and $m=m(\varepsilon,\delta)$, we have
@@
In our setup we emphasize time invariance, i.e. that the same base function $f\in{\mathcal{F}}$ is used in every step of autoregressive generation (e.g., a transformer with the same parameters, in a practical setting of interest). We then need to learn just a single function $f\in{\mathcal{F}}$, and our goal is for the learning complexity to be independent of (or at least nearly independent of) $T$. Time-invariance here can also be viewed as “parameter sharing” (Rajaraman et al., 2021) across steps of generation: for a parametric class =${\mathcal{F}}=\left\{f_{\bm{\theta}}:{\bm{\theta}}\in{\bm{\Theta}}\right\}$, the parameter ${\bm{\theta}}$ is shared among all time steps and we only need to learn a single ${\bm{\theta}}$. This invariance is in contradistinction to a Time-Dependent notion of autoregressive learning suggested by Malach (2023), where a different base predictor $f_{t}\in{\mathcal{F}}$ is used in every step $t$, yielding the Time-Dependent Chain-of-Thought class
@@
Malach considered learning the class ${\mathcal{F}}^{\textnormal{TD-}{\sf CoT}\textnormal{-}{T}}$, i.e. with the CoT available during training (as in our Definition 2) [^5] In Malach (2023), the learning goal is to be able to generate the entire CoT, not just the final answer. This is different from our Definition 2, but this is not a substantial difference.. One can also consider end-to-end learning (as in our Definition 1) in such a time-dependent setting, which amounts to PAC-Learning the hypothesis class:
@@
Sample complexities obtained by Malach all scale linearly with $T$, and this is unavoidable without time-invariance—see also discussion in Section 3. Our goal is to avoid such dependence.
@@
Although all our learning guarantees are for learning over the entire ${\mathcal{X}}=\Sigma^{*}$, and do not have any dependence on the length of the input, when discussing computational complexity, the length of the input will be important. Likewise, instead of considering a fixed base class ${\mathcal{F}}$, we must consider a family ${\mathcal{F}}_{d}$ parametrized by (one or more) size parameters $d$. We will say that ${\mathcal{F}}_{d}^{{\sf e2e}\textnormal{-}{T}}$ is ${\sf e2e}$ or ${\sf CoT}$ learnable in $\mathsf{time}(n,d,T,\varepsilon,\delta)$ using some learning algorithm [^6] Formally, to allow uniform algorithms, we can think of $A$ being implicitly passed $T$ and $d$. $A$, if over the domain ${\mathcal{X}}=\Sigma^{\leq n}$ (i.e. restricted to prompts of length at most $n$), $A(S)$ runs in time at most $\mathsf{time}(n,T,d,\varepsilon,\delta)$ almost surely. For an expression $\kappa(\psi_{1},\dots,\psi_{k})$ in scalar quantities $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, we say that $\kappa$ is in ${\mathcal{P}oly}(\psi_{1},\dots,\psi_{k})$ if $\kappa$ is uniformly bounded by a polynomial for all $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, i.e. there exists a polynomial $p:\mathbb{R}^{k}\to\mathbb{R}$ such that for every $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, we have $\kappa(\psi_{1},\dots,\psi_{k})\leq p(\psi_{1},\dots,\psi_{k})$.
@@ 3 Statistical and Computational Complexity for a General ${\mathcal{F}}$
In this section, we discuss the learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ for a hypothesis class ${\mathcal{F}}$ in terms of general properties of ${\mathcal{F}}$; see Table 1. Complete proofs can be found in Appendix B.
@@
(Theorem 3.5) (dominated by $\mathrm{VCdim}({\mathcal{F}})\log T$)
Computational complexity when Not necessarily ${\mathcal{P}oly}(n,T,d)$ ${\mathcal{P}oly}(n,T,d)$
Cons for $({\mathcal{F}}_{d})$ is tractable (Section 4) (Corollary 3.7)
Table 1: A summary of the results in Section 3 for time-invariant autoregressive learning for a general base class ${\mathcal{F}}$. Each sample complexity row indicates the best possible guarantee, up to $O(\log(1/(\varepsilon\delta))/\varepsilon)$, on the sample complexity of ${\sf e2e}$ or ${\sf CoT}$ learning under the corresponding assumption on ${\mathcal{F}}$. All indicated sample complexities are tight except for $\log T$ factors, i.e. for each cell there exists classes ${\mathcal{F}}$ matching the indicated upper bound except for the $\log T$ factor (see Appendix E).
@@ 3.1 Learnability and Sample Complexity
While this is a promising start, it is natural to wonder if analogous results hold for more general function classes satisfying less stringent complexity bounds, such as requiring only bounded VC dimension. Indeed, for classes over $\Sigma=\{0,1\}$ (we will return to general alphabets in Remark 3.1 at the end of the section) we can bound $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=O(T\cdot\mathrm{VCdim}({\mathcal{F}}))$, which leads to the following sample complexity guarantee for ${\sf e2e}$-learning.
@@
The linear dependence on $T$ is extremely disappointing, as our goal with time-invariance is to avoid such generation-length dependence. In fact, we can learn the time-dependent ${\mathcal{F}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$, with $O(T\cdot\mathrm{VCdim}({\mathcal{F}})\log T)$ samples, i.e. as in Theorem 3.2 up to a $\log T$ factor (Theorem E.2 in Section E.2). Unfortunately, even with time-invariance, the linear dependence on $T$ is unavoidable:
@@
For every $D,T\in{\mathbb{N}}_{+}$, there exists a base class ${\mathcal{F}}\subseteq\{0,1\}^{\{0,1\}^{*}}$ with $\mathrm{VCdim}({\mathcal{F}})=D$, and a distribution ${\mathcal{D}}$ over $\{0,1\}^{n}$ for $n=\lceil\log(DT)\rceil+1$ such that for any learning rule $A$ there exists $f_{*}\in{\mathcal{F}}$, s.t. with probability at least $0.8$ over $S_{{\sf e2e}}$ of size $m<\frac{DT}{2}$ (sampled as in Definition 1), we have $L_{{\mathcal{D}},f_{*}}^{\textnormal{0-1}}(A(S_{{\sf e2e}}))\geq\frac{1}{4}$.
@@
That is, while the VC dimension of the base class ${\mathcal{F}}$ is not sufficient for ensuring non-invariant sample complexity with end-to-end learning, having the chain-of-thought available during training does allow for a (nearly) generation-length independent sample complexity based only on the VC dimension of the base class—see Table 1.
@@
In Appendix B, Corollaries B.2, B.4 and B.6, we provide extensions of our Theorems 3.2, 3.5 and 3.4 for any general finite alphabet $\Sigma$. The bounds are similar, with the VC dimension replaced with the Natarajan dimension and the Littlestone dimension replaced with the sequential shattering dimension, but there is also an additional factor of $\log\left\lvert{\Sigma}\right\rvert$ in the sample complexities. We also show in Theorem E.5 that with the infinite alphabet $\Sigma=\mathbb{R}$, even if the base class ${\mathcal{F}}$ has bounded VC-subgraph dimension (a.k.a. Pollard, or pseudo dimension), the end-to-end class ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ could have infinite VC-subgraph dimension.
@@
CoT Learnability in terms of VC Dimension (Theorem 3.4).
@@
$\Gamma_{{\mathcal{L}}^{{\sf CoT}\textnormal{-}{T}}}(m)\leq\Gamma_{{\mathcal{F}}}(mT)\leq\left(\frac{e\,m\,T}{\scriptstyle{\mathrm{VCdim}({\mathcal{F}})}}\right)^{\mathrm{VCdim}({\mathcal{F}})}\Rightarrow\mathrm{VCdim}({\mathcal{L}}^{{\sf CoT}\textnormal{-}{T}})=O(\mathrm{VCdim}({\mathcal{F}})\log T)$
@@
e2e Learnability in terms of VC Dimension (Theorem 3.2).
@@
$\Gamma_{{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}}(m)\leq\Gamma_{{\mathcal{F}}}(m\cdot 2^{T})\leq\left(\frac{e\,m\,2^{T}}{\scriptstyle{\mathrm{VCdim}({\mathcal{F}})}}\right)^{\mathrm{VCdim}({\mathcal{F}})}\Rightarrow\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=O(T\,\mathrm{VCdim}({\mathcal{F}})).$
e2e Learnability in terms of the Littlestone Dimension (Theorem 3.5).
@@
$\Gamma_{{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}}(m)\leq\left(\#\text{ behaviors on a binary tree of depth $m(T+1)$}\right)\leq\left(\frac{2em(T+1)}{\operatorname{Ldim}({\mathcal{F}})}\right)^{\operatorname{Ldim}({\mathcal{F}})},$

### v2

@@ 1 Introduction
In our view, a central component of autoregressive generation is that it is time-invariant, i.e., the same next-token-generator, with the identical parameters (e.g. a transformer with the same weights) is used at each step of generation, to generate each token as a function of the prefix thus far. Throughout the paper we emphasize how such time-invariance is crucial for allowing learning with sample complexity independent (perhaps up to a log-factor) of the generation length, and later on, of the compute time (number of steps) of a learned process. In this crucial regard, we deviate significantly from another recent (and inspiring) attempt to formalize autoregressive learning by Malach (2023), who considered time-dependent autoregressive learning. We directly contrast with Malach’s approach as we progress through our presentation.
We begin by presenting a formal setting for time-invariant autoregressive Chain-of-Thought generation and learning using an abstract base class of next-token-generators (Section˜2), studying the basics of learning in this framework for general and abstract base classes (Section˜3), and, as an example, for what is perhaps the simplest possible base class, namely linear thresholds (Section˜4). We consider training both with and without explicit Chain-of-Thought supervision. Already here we can start seeing the computational benefits of Chain-of-Thought learning and the statistical benefits of time-invariant autoregressive generation. We then turn to the more ambitious goal of “universal” learning of computable functions and see how far the benefits of time-invariance and autoregressive Chain-of-Thought take us, and how attention arises naturally (Sections˜5 and 6).
@@ 2 Time-Invariant Autoregressive Chain-of-Thought Learning
For a finite set of tokens $\Sigma$, a next token generator is a mapping $f:\Sigma^{*}\to\Sigma$. On input ${\bm{x}}\in\Sigma^{*}$, autoregressive Chain-of-Thought (CoT) generation will start with a string containing only the input and will repeatedly apply $f$ to the string and append the newly generated token to it (see Figure˜1). Formally, for a next-token generator $f$, we define the apply-and-append mapping $\overline{f}({\bm{x}}):{\bm{x}}\mapsto\textnormal{append}({\bm{x}},f({\bm{x}}))\in(\Sigma^{\left\lvert{{\bm{x}}}\right\rvert+1}$), [^2] For any ${\bm{x}}\in\Sigma^{*}$, we use $|{\bm{x}}|$ to denote its length. We use negative indices to refer to elements of a vector or string from the end, i.e. ${\bm{v}}[-1]$ is the last element. We use inclusive slice indexing, so that ${\bm{v}}[i:j]$ are the elements from position $i$ to $j$ inclusive. Dropping $i$ or $j$ means that the slice extends from the beginning or to the end, respectively. and apply $\overline{f}$ iteratively for $T$ steps to obtain a mapping from $\bm{x}$} to its $T$-step Chain-of-Thought:
@@
For any base class ${\mathcal{F}}=\{f:\Sigma^{*}\rightarrow\Sigma\}\subseteq\Sigma^{\Sigma^{*}}$ of generators, and a generation length $T$, we define the corresponding end-to-end and CoT classes:
@@
For simplicity, throughout the paper we consider a fixed thought generation length (i.e. number of iterations) $T$. This number of iterations $T$ will be a crucial parameter for us and we will study how the learning complexity and expressivity depend on TT.333 One could also define a variant with variable thought lengths with a special token indicating when to stop. But if we would anyway be interested in the maximal thought length, we could just as well extend each thought with a padding token to make them all of the same length.
@@
For a base class ${\mathcal{F}}\subseteq\Sigma^{\Sigma^{*}}$, and generation length $T\in{\mathbb{N}}_{+}$, we say that ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf e2e$}-learnable over a domain ${\mathcal{X}}\subseteq\Sigma^{*}$, with sample complexity $m(\varepsilon,\delta)$ if there exists a learning rule $A:({\mathcal{X}}\times\Sigma)^{*}\to\Sigma^{{\mathcal{X}}}$ such that for every distribution $\mathcal{D}$} over $\mathcal{X}$} and $f_{*}\in{\mathcal{F}}$, for every $0<\varepsilon,\delta<1$, with probability at least $1-\delta$ over ${\bm{x}}_{1},\dots,{\bm{x}}_{m}\sim_{iid}{\mathcal{D}}$, with $S_{{\sf e2e}}=(({\bm{x}}_{1},y_{1}),\dots,({\bm{x}}_{m},y_{m}))$, $y_{i}=f_{*}^{{\sf e2e}\textnormal{-}{T}}({\bm{x}}_{i})$ and $m=m(\varepsilon,\delta)$, we have $L_{{\mathcal{D}},f_{*}}^{\textnormal{0-1}}(A(S_{{\sf e2e}}))\leq\varepsilon$.
Definition˜1 amounts precisely to the standard definition of PAC-learning the hypothesis class ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$, but we give it explicitly in order to contrast it with learning in situations where the entire Chain-of-Thought is available during training. In contemporary LLM training, such CoT supervision can be found in data sources such scrapped over internet consisting of step-by-step answers to questions in discussion forums and in textbooks (Lewkowycz et al., 2022, Li et al., 2023, Gao et al., 2020, Achiam et al., 2023), as well as in training data specifically curated (Ott et al., 2023, Chung et al., 2024, Hendrycks et al., 2021).
@@
For a base class ${\mathcal{F}}\subseteq\Sigma^{\Sigma^{*}}$ and generation length $T\in{\mathbb{N}}_{+}$, we say that [^4] We emphasize that $\sf CoT$}-learnability is a property of the base class $\mathcal{F}$} and length $T$, not only of the set ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$. ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf CoT$}-learnable over a domain ${\mathcal{X}}\subseteq\Sigma^{*}$, with sample complexity $m(\varepsilon,\delta)$, if there exists a learning rule $A:({\mathcal{X}}\times\Sigma^{T})^{*}\to\Sigma^{{\mathcal{X}}}$ such that for every distribution $\mathcal{D}$} over $\mathcal{X}$} and $f_{*}\in{\mathcal{F}}$, for every $0<\varepsilon,\delta<1$, with probability at least $1-\delta$ over ${\bm{x}}_{1},\dots,{\bm{x}}_{m}\sim_{iid}{\mathcal{D}}$, with $S_{{\sf CoT}}=({\bm{z}}_{1},\dots,{\bm{z}}_{m})$ where ${\bm{z}}_{i}=f_{*}^{{\sf CoT}\textnormal{-}{T}}({\bm{x}}_{i})$, and $m=m(\varepsilon,\delta)$, we have
@@
In our setup we emphasize time invariance, i.e. that the same base function $f\in{\mathcal{F}}$ is used in every step of autoregressive generation (e.g., a transformer with the same parameters, in a practical setting of interest). We then need to learn just a single function $f\in{\mathcal{F}}$, and our goal is for the learning complexity to be independent of (or at least nearly independent of) $T$. Time-invariance here can also be viewed as “parameter sharing” (Rajaraman et al., 2021) across steps of generation: for a parametric class ${\mathcal{F}}=\left\{f_{\bm{\theta}}:{\bm{\theta}}\in{\bm{\Theta}}\right\}$, the parameter $\bm{\theta}$} is shared among all time steps and we only need to learn a single $\bm{\theta}$}. This invariance is in contradistinction to a Time-Dependent notion of autoregressive learning suggested by Malach (2023), where a different base predictor $f_{t}\in{\mathcal{F}}$ is used in every step $t$, yielding the Time-Dependent Chain-of-Thought class
@@
Malach considered learning the class ${\mathcal{F}}^{\textnormal{TD-}{\sf CoT}\textnormal{-}{T}}$, i.e. with the CoT available during training (as in our Definition˜2) [^5] In Malach (2023), the learning goal is to be able to generate the entire CoT, not just the final answer. This is different from our Definition 2, but this is not a substantial difference.. One can also consider end-to-end learning (as in our Definition˜1) in such a time-dependent setting, which amounts to PAC-Learning the hypothesis class:
@@
Sample complexities obtained by Malach all scale linearly with $T$, and this is unavoidable without time-invariance—see also discussion in Section˜3. Our goal is to avoid such dependence.
@@
Although all our learning guarantees are for learning over the entire ${\mathcal{X}}=\Sigma^{*}$, and do not have any dependence on the length of the input, when discussing computational complexity, the length of the input will be important. Likewise, instead of considering a fixed base class $\mathcal{F}$}, we must consider a family ${\mathcal{F}}_{d}$ parametrized by (one or more) size parameters $d$. We will say that ${\mathcal{F}}_{d}^{{\sf e2e}\textnormal{-}{T}}$ is $\sf e2e$} or $\sf CoT$} learnable in $\mathsf{time}(n,d,T,\varepsilon,\delta)$ using some learning algorithm [^6] Formally, to allow uniform algorithms, we can think of $A$ being implicitly passed $T$ and $d$. AA, if over the domain ${\mathcal{X}}=\Sigma^{\leq n}$ (i.e. restricted to prompts of length at most $n$), $A(S)$ runs in time at most $\mathsf{time}(n,T,d,\varepsilon,\delta)$ almost surely. For an expression $\kappa(\psi_{1},\dots,\psi_{k})$ in scalar quantities $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, we say that $\kappa$ is in ${\mathcal{P}\hskip-1.0ptoly}(\psi_{1},\dots,\psi_{k})$ if $\kappa$ is uniformly bounded by a polynomial for all $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, i.e. there exists a polynomial $p:\mathbb{R}^{k}\to\mathbb{R}$ such that for every $(\psi_{1},\dots,\psi_{k})\in\mathbb{R}^{k}$, we have $\kappa(\psi_{1},\dots,\psi_{k})\leq p(\psi_{1},\dots,\psi_{k})$.
@@ 3 Statistical and Computational Complexity for a General $\mathcal{F}$}
In this section, we discuss the learnability of ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ for a hypothesis class $\mathcal{F}$} in terms of general properties of $\mathcal{F}$}; see Table˜1. Complete proofs can be found in Appendix˜B.
@@
(Theorem˜3.5) (dominated by $\mathrm{VCdim}({\mathcal{F}})\log T$)
Computational complexity when Not necessarily ${\mathcal{P}\hskip-1.0ptoly}(n,T,d)$ ${\mathcal{P}\hskip-1.0ptoly}(n,T,d)$
Cons for $({\mathcal{F}}_{d})$ is tractable (Section˜4) (Corollary˜3.7)
Table 1: A summary of the results in Section˜3 for time-invariant autoregressive learning for a general base class $\mathcal{F}$}. Each sample complexity row indicates the best possible guarantee, up to $O(\log(1/(\varepsilon\delta))/\varepsilon)$, on the sample complexity of $\sf e2e$} or $\sf CoT$} learning under the corresponding assumption on $\mathcal{F}$}. All indicated sample complexities are tight except for $\log T$ factors, i.e. for each cell there exists classes $\mathcal{F}$} matching the indicated upper bound except for the $\log T$ factor (see Appendix˜E).
@@ 3.1 Learnability and Sample Complexity
While this is a promising start, it is natural to wonder if analogous results hold for more general function classes satisfying less stringent complexity bounds, such as requiring only bounded VC dimension. Indeed, for classes over $\Sigma=\{0,1\}$ (we will return to general alphabets in Remark˜3.1 at the end of the section) we can bound $\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=O(T\cdot\mathrm{VCdim}({\mathcal{F}}))$, which leads to the following sample complexity guarantee for $\sf e2e$}-learning.
@@
The linear dependence on $T$ is extremely disappointing, as our goal with time-invariance is to avoid such generation-length dependence. In fact, we can learn the time-dependent ${\mathcal{F}}^{\textnormal{TD-}{\sf e2e}\textnormal{-}{T}}$, with $O(T\cdot\mathrm{VCdim}({\mathcal{F}})\log T)$ samples, i.e. as in Theorem˜3.2 up to a $\log T$ factor (Theorem˜E.2 in Section˜E.2). Unfortunately, even with time-invariance, the linear dependence on $T$ is unavoidable:
@@
For every $D,T\in{\mathbb{N}}_{+}$, there exists a base class ${\mathcal{F}}\subseteq\{0,1\}^{\{0,1\}^{*}}$ with $\mathrm{VCdim}({\mathcal{F}})=D$, and a distribution $\mathcal{D}$} over $\{0,1\}^{n}$ for $n=\lceil\log(DT)\rceil+1$ such that for any learning rule $A$ there exists $f_{*}\in{\mathcal{F}}$, s.t. with probability at least $0.8$ over $S_{{\sf e2e}}$ of size $m<\frac{DT}{2}$ (sampled as in Definition˜1), we have $L_{{\mathcal{D}},f_{*}}^{\textnormal{0-1}}(A(S_{{\sf e2e}}))\geq\frac{1}{4}$.
@@
As $\textsc{Cons}_{{\sf CoT}}$ is a special case of [^7] As with many standard learning rules, the rule is not a specific mapping, but a specification that can be implemented by many different mappings. Since any $\sf CoT$} consistent $f$ is also $\sf e2e$} consistent, any valid response of $\textsc{Cons}_{{\sf CoT}}$ is also valid for $\textsc{Cons}_{{\sf e2e}}$, but not vice versa. $\textsc{Cons}_{{\sf e2e}}$, we have that $\textsc{Cons}_{{\sf CoT}}$ enjoys the same guarantees as $\textsc{Cons}_{{\sf e2e}}$, including the cardinality based Theorem˜3.1. But for VC base classes, $\textsc{Cons}_{{\sf CoT}}$ enjoys a much stronger (nearly independent of $T$) guarantee than in Theorem˜3.2:
@@
That is, while the VC dimension of the base class $\mathcal{F}$} is not sufficient for ensuring $T$ independent sample complexity for end-to-end learning, having the chain-of-thought available during training does allow for a (nearly) generation-length independent sample complexity based only on the VC dimension of the base class—see Table˜1.
@@
Our Theorems˜3.2, 3.4 and 3.5 follow from simple but elegant covering number arguments (sketched out in Section˜3.2), followed by the standard uniform convergence of the empirical and population losses.
@@
In Appendix˜B, Corollaries˜B.2, B.4 and B.6, we provide extensions of our Theorems˜3.2, 3.5 and 3.4 for any general finite alphabet $\Sigma$. The bounds are similar, with the VC dimension replaced with the Natarajan dimension and the Littlestone dimension replaced with the sequential shattering dimension, but there is also an additional factor of $\log\left\lvert{\Sigma}\right\rvert$ in the sample complexities. We also show in Theorem˜E.5 that with the infinite alphabet $\Sigma=\mathbb{R}$, even if the base class $\mathcal{F}$} has bounded VC-subgraph dimension (a.k.a. Pollard, or pseudo dimension), the end-to-end class ${\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}$ could have infinite VC-subgraph dimension.
@@
CoT Learnability in terms of VC Dimension (Theorem˜3.4).
@@
$\Gamma_{{\mathcal{L}}^{{\sf CoT}\textnormal{-}{T}}}(m)\leq\Gamma_{{\mathcal{F}}}(mT)\leq\left(\frac{e\,m\,T}{\scriptstyle{\mathrm{VCdim}({\mathcal{F}})}}\right)^{\mathrm{VCdim}({\mathcal{F}})}\hskip-5.69054pt\Rightarrow\mathrm{VCdim}({\mathcal{L}}^{{\sf CoT}\textnormal{-}{T}})=O(\mathrm{VCdim}({\mathcal{F}})\log T)$
@@
e2e Learnability in terms of VC Dimension (Theorem˜3.2).
@@
$\Gamma_{{\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}}}(m)\leq\Gamma_{{\mathcal{F}}}(m\cdot 2^{T})\leq\left(\frac{e\,m\,2^{T}}{\scriptstyle{\mathrm{VCdim}({\mathcal{F}})}}\right)^{\mathrm{VCdim}({\mathcal{F}})}\hskip-8.53581pt\Rightarrow\mathrm{VCdim}({\mathcal{F}}^{{\sf e2e}\textnormal{-}{T}})=O(T\,\mathrm{VCdim}({\mathcal{F}})).$
e2e Learnability in terms of the Littlestone Dimension (Theorem˜3.5).
@@
[…]